# Mapping bipartite networks into multidimensional hyperbolic spaces

### v1

@@ Abstract
Bipartite networks appear in many real-world contexts, linking entities across two distinct sets. They are often analyzed via one-mode projections, but such projections can introduce artificial correlations and inflated clustering, obscuring the true underlying structure. In this paper, we propose a geometric model for bipartite networks that leverages the high levels of bipartite four-cycles as a measure of clustering to place both node types in the same similarity space, where link probabilities decrease with distance. Additionally, we introduce B-Mercator, an algorithm that infers node positions from the bipartite structure. We evaluate its performance on diverse datasets, illustrating how the resulting embeddings improve downstream tasks such as node classification and distance-based link prediction in machine learning. These hyperbolic embeddings also enable the generation of synthetic networks with node features closely resembling real-world ones, thereby safeguarding sensitive information while allowing secure data sharing. In addition, we show how preserving bipartite structure avoids the pitfalls of projection-based techniques, yielding more accurate descriptions and better performance. Our method provides a robust framework for uncovering hidden geometry in complex bipartite systems.
I Introduction
Bipartite networks lie at the heart of countless real-world applications, linking authors to the articles they write Newman (2001a, b, c, 2004), users to the products they consume Koren et al. (2009); Konstan et al. (1997); Linden et al. (2003), people to the groups they belong to Davis et al. (2009), or countries to the languages they speak Consortium (2024). They also arise naturally in metabolic networks, where metabolites are connected to the chemical reactions or enzymes that transform them Jeong et al. (2000); Ravasz et al. (2002); Serrano et al. (2012), in plant-pollinator networks Bascompte et al. (2003); Olesen et al. (2007), and in machine learning applications, where nodes have associated features used to feed graph neural networks Aliakbarisani et al. (2023). By design, each bipartite system splits its nodes into two disjoint sets, with no edges connecting nodes within the same set. This seemingly simple rule nonetheless yields rich and complex connectivity patterns, enabling researchers to model collaboration, consumption, and association processes across diverse domains. Yet bipartite networks have historically garnered less attention than their unipartite counterparts. Given their ubiquity and straightforward interpretability, a renewed focus on bipartite structures is both timely and necessary to fully capture the multiple facets of interaction present in many real-world complex systems and data structures.
@@ I Introduction
To overcome these limitations and more accurately capture the true structure of bipartite systems, it is critical to develop a modeling framework that treats bipartite networks directly rather than relying solely on their one-mode projections. In this paper, we propose to use network geometry to make sense of bipartite networks and to find geometric representations for this class of systems. In recent years, network geometry Boguñá et al. (2021) has been extremely successful in explaining undirected Serrano et al. (2008), directed Allard et al. (2024), weighted Allard et al. (2017), and multiplex networks Kleineberg et al. (2016, 2017) in many real systems and has also been extended to bipartite settings, proposing that nodes of both types can indeed lie in a shared latent space with connection probabilities governed by their mutual distances Serrano et al. (2012); Kitsak et al. (2017). Despite the fact that bipartite networks do not contain triangles by definition (the signature of any metric space), it is still possible to define a clustering coefficient by counting cycles of length four (i.e., squares). Empirical analyses have shown that these four-cycles can be abundant in real bipartite networks, leading to high effective clustering values Opsahl (2013); Aliakbarisani et al. (2023). This finding suggests that, indeed, bipartite networks may be embedded in a metric space where the likelihood of a connection between two nodes of different types decreases with the distance separating them.
@@
Assuming that a bipartite network can be embedded in a similarity (or metric) space, a logical step is to seek methods for inferring the positions of its nodes within this space from real data. Previous approaches typically rely on one-mode projections of bipartite networks –using, for instance, the D-Mercator tool García-Pérez et al. (2019); Jankowski et al. (2023)– to embed each type of node. However, such projections often introduce artificial correlations and inflated clustering, reducing embedding accuracy. Here, following results in Serrano et al. (2012); Kitsak et al. (2017), we introduce the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model in which nodes of both types lie in the same $D$-dimensional similarity space. Using this model, we propose B-Mercator, an algorithm designed specifically for bipartite networks that enables the creation of multidimensional hyperbolic maps of real bipartite datasets. To illustrate its capabilities, we embeded and analyzed three datasets: Unicodelang Consortium (2024), which links countries to the languages spoken in those regions, Metabolic King et al. (2016), in which metabolites are connected through chemical reactions, and Flavor Ahn et al. (2011), which connects food ingredients to their corresponding flavor compounds. Furthermore, we show how B-Mercator can be applied to supervised machine learning tasks, including node classification and link prediction, yielding a significant performance gain with respect to state of the art methods, especially when a strong correlation exists between nodes’ labels and their feature distributions.
@@
Figure 1: Schematic representation of the bipartite-$\mathbb{S}^{D}$ model in (a) $D=1$ and (b) $D=2$. Nodes A are shown as circles whereas nodes B are shown as squares whose sizes are proportional to the nodes’ expected degrees. The angular distances between nodes A and B are highlighted. Light grey lines represent the edges in the bipartite network generated by Eq. (1).
II Results
II.1 The bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model and B-Mercator
@@
Given a bipartite network whose links connect type-A and type-B nodes, B-Mercator finds an embedding of both node types on the surface of a $D$-sphere $\mathbb{S}^{D}$ (or, equivalently, within a $(D+1)$-sphere in the hyperbolic space $\mathbb{H}^{D+1}$), which serves as a likely realization of the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model. In the hyperbolic representation, the coordinates of nodes on the $D$-sphere represent their positions in the similarity space, whereas the radial coordinate within the hyperbolic $(D+1)$-sphere encodes their popularity, as reflected by their expected degree. Figure 1 shows two examples of bipartite networks generated by the bipartite-$\mathbb{S}^{D}$ model for $D=1$ and $D=2$. In a nutshell, pairs of type-A and -B nodes are connected with a probability that depends on their distance on the sphere, rescaled by the product of their expected degrees (represented by the size of the nodes in the figure). In the hyperbolic representation, each node is assigned a radial coordinate so that the connection probability becomes a function of the hyperbolic distance between the two nodes, thereby making the model a random geometric graph in hyperbolic geometry. In both representations, the model includes an inverse temperature $\beta_{b}$, which modulates the amount of noise in the system and thereby gauges the coupling between the network topology and its geometric characteristics. A full description of the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model and the technical details of the B-Mercator embedding algorithm are provided in the Methods sections IV.1 and IV.2.
II.2 Validation
@@
To validate our method, we generated synthetic bipartite networks from the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model and measured the quality of the estimated node positions. Figure 2 shows comparisons between the true and inferred coordinates for type A and B nodes. One can observe a high Pearson correlation coefficient for $D=1$ and $D=2$, corroborating the effectiveness of our embedding technique. For more examples in $D=1$ and $D=2$ as well as $D=3$, see Supplementary Figures 2-4. B-Mercator can also infer the inverse temperature $\beta_{b}$ as shown in Supplementary Figure 5. Moreover, we tested the reproducibility of the topological properties of the original network by generating an ensemble of synthetic networks using the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model with the inferred parameters and nodes’ coordinates (see Supplementary Figures 6-10). The degree distributions and clustering spectra of type A and B nodes were very well reproduced. We also observed a good agreement between the empirical and theoretical connection probabilities. These findings confirm that B-Mercator stands out as a high-quality algorithm, accurately reconstructing the nodes’ coordinates of synthetic networks. Moreover, it reliably determines all other model parameters, such as hidden degrees and inverse temperature, regardless of the network’s dimensionality.
@@
Figure 2: Validation of B-Mercator on synthetic bipartite networks. Relationship between the original and the inferred coordinates of the (a, b) bipartite-$\mathbb{S}^{1}$ and (c, d) bipartite-$\mathbb{S}^{2}$ models. In the top left corner of each figure, we report the value of the Pearson correlation coefficient between the inferred and original coordinates. Since the inferred coordinates for $D=2$ might be rotated, we transform them to minimize the average angular distance between the original and inferred coordinates (cf. Supplementary Section III in Jankowski et al. (2023)). Parameters for (a, b) ($D,N_{A},N_{B},\gamma_{A},\gamma_{B},\langle k_{A}\rangle,\beta_{b}$) = ($1,500,1000,2.7,2.1,10,1.5$), for (c, d) ($D,N_{A},N_{B},\gamma_{A},\gamma_{B},\langle k_{A}\rangle,\beta_{b}$) = ($2,500,1000,2.7,2.7,10,6$).
II.3 Bipartite greedy routing
@@
In order to establish a meaningful geometric representation of a bipartite network at the global scale, we introduce bipartite greedy routing (BGR) as a practical tool to infer the network’s effective dimension Boguñá et al. (2009). The idea is to first embed the bipartite network into a latent geometric space with B-Mercator and then test how well nodes can route information by simply forwarding messages to their neighbors closest to the destination in that space. By systematically evaluating the success of these greedy routes –measured, for instance, by the probability that messages reach their targets without getting stuck– we gain insights into the dimensional structure underlying the network. A lower-dimensional latent space often requires fewer “hops” and exhibits more consistent greedy paths, thus reflecting more coherent topological structures. Conversely, if greedy routing frequently fails or requires excessive detours, it suggests a higher-dimensional or more complex geometry. In this way, the performance of greedy routing serves as an indicator of how well the bipartite graph can be embedded in a space of a given dimension, effectively allowing us to determine the dimension that best captures its structure.
We implemented a BGR protocol in which both the origin and destination can be either type A or type B nodes thus defining four variants. In Figure 3a, we depict a schematic picture of the BGR for the A–B variant, i.e., when the source node is a type A node ($S_{A}$), and the target node is a type B node ($T_{B}$). The message is forwarded from $S_{A}$ to the type B node that is hyperbolically closest to the target. Since the node $B_{1}$ is not the destination, the message is forwarded again to a type A node. The process is repeated until the destination is reached or the message becomes stuck. Then, the BGR protocol is executed for a large number of randomly chosen node pairs to assess the global network’s geometric properties.
We tested the BGR protocol in synthetic networks generated from the bipartite-$\mathbb{S}^{D}$ model. First, we generated networks with specific dimensionality and topological properties, and we obtained their hyperbolic maps by embedding them using B-Mercator with different embedding dimensions. The performance of BGR was assessed based on two key measures: the proportion of messages that successfully reach their destination $p_{s}$, i.e., the success rate, and the mean stretch, where stretch is defined for each path connecting a source and target node as the ratio of the hop count of a successful greedy path to the shortest path. In Figure 3b, we show the success rate as a function of the embedding dimension for networks generated using the bipartite-$\mathbb{S}^{D}$ model with dimensions ranging from $D=1$ to $D=4$. Interestingly, the performance of BGR is optimal (in terms of $p_{s}$ and mean stretch, see Supplementary Figure 12) when B-Mercator is used with the same dimension that was used to generate the network, thus justifying BGR as an alternative method to infer the effective dimension of real networks, different from the topological-based method introduced in Almagro et al. (2022). These results are consistent across all variants of BGR (see Supplementary Figure 11) and corroborate the results obtained for unipartite networks Jankowski et al. (2023).
@@
Figure 3: Bipartite greedy routing in synthetic networks. (a) Schematic view of the greedy routing protocol. We select a type A node as an origin ($\mathbf{S_{A}}$) and a type B ($\mathbf{T_{B}}$) as a destination. The red arrows show how the package/message is forwarded towards the destination. In the second example, we select two type B nodes as source ($\mathbf{S_{B^{\prime}}}$) and destination ($\mathbf{T_{B^{\prime\prime}}}$) and outline the greedy path with purple color. The line width is proportional to the connection probability (Eq. 1). (b) Success rate as a function of embedded dimension for networks generated in $D=\{1,2,3,4\}$. We consider here a navigation protocol where a source is a type A node and a destination a type B one. Results are obtained by averaging over 10 realizations with $(N_{A},N_{B},\gamma_{A},\gamma_{B},\left<k\right>,\beta_{b})$ = $(500,500,2.5,2.5,10,1.5)$.
II.4 Embedding of real bipartite networks

### v2

@@ Abstract
Bipartite networks appear in many real-world contexts, linking entities across two distinct sets. One-mode projections, though widely used, can introduce artificial correlations and inflated clustering, obscuring the true underlying structure. In this work, we propose a geometric model for bipartite networks that leverages the high levels of bipartite four-cycles as a measure of clustering and embeds both node types into a shared similarity space, where the link probability decreases with distance. Additionally, we introduce B-Mercator, an algorithm that infers node positions from the bipartite structure. We evaluate its performance on diverse datasets, illustrating how the resulting embeddings improve downstream tasks such as node classification and distance-based link prediction. These hyperbolic embeddings enable realistic synthetic network generation with node features mirroring real data. By preserving the bipartite structure, our approach avoids projection biases, offering more accurate structural descriptions and providing a robust framework for uncovering hidden geometry in bipartite networks.
Bipartite networks lie at the heart of countless real-world applications, linking authors to the articles they write Newman (2001a, b, c, 2004), users to the products they consume Koren et al. (2009); Konstan et al. (1997); Linden et al. (2003), people to the groups they belong to Davis et al. (2009), or countries to the languages they speak Unicode Consortium (2024). They also arise naturally in metabolic networks, where metabolites are connected to the chemical reactions or enzymes that transform them Jeong et al. (2000); Ravasz et al. (2002); Serrano et al. (2012), in plant-pollinator networks Bascompte et al. (2003); Olesen et al. (2007), and in machine learning applications, where nodes have associated features used to feed graph neural networks Aliakbarisani et al. (2025). By definition, each bipartite system splits its nodes into two disjoint sets, with no edges connecting nodes within the same set. This seemingly simple rule nonetheless yields rich and complex connectivity patterns, enabling researchers to model collaboration, consumption, and association processes across diverse domains. Yet bipartite networks have historically garnered less attention than their unipartite counterparts. Given their ubiquity and straightforward interpretability, a renewed focus on bipartite structures is both timely and necessary to fully capture the multiple facets of interaction present in many real-world complex systems and data structures.
@@
To overcome these limitations and more accurately capture the true structure of bipartite systems, it is critical to develop a modeling framework that treats bipartite networks directly rather than relying solely on their one-mode projections. Thus, we turn our attention to the latent space models, initially developed for social networks Hoff et al. (2002) and extended to temporal bipartite networks Friel et al. (2016) and to growing bipartite generative models Filho and O’Neale (2020). Within this family of models, we propose to use network geometry to make sense of bipartite networks and to find hyperbolic geometric representations for this class of systems Krioukov et al. (2010). Hyperbolic geometry is particularly useful for capturing the hierarchical organization of complex networks Ortiz et al. (2020). In recent years, network geometry Boguñá et al. (2021) has been extremely successful in explaining undirected Serrano et al. (2008), directed Allard et al. (2024), weighted Allard et al. (2017), and multiplex networks Kleineberg et al. (2016, 2017) in many real systems and has also been extended to bipartite settings, proposing that nodes of both types can indeed lie in a shared latent space with connection probabilities governed by their mutual distances Serrano et al. (2012); Kitsak et al. (2017). Despite the fact that bipartite networks do not contain triangles by definition (the signature of any metric space), it is still possible to define a clustering coefficient by counting cycles of length four (i.e., squares). Empirical analyses have shown that these four-cycles can be abundant in real bipartite networks, leading to high effective clustering values Opsahl (2013); Aliakbarisani et al. (2025). This finding suggests that, indeed, bipartite networks may be embedded in a metric space where the likelihood of a connection between two nodes of different types decreases with the distance separating them.
Assuming that a bipartite network can be embedded in a similarity (or metric) space, a logical step is to seek methods for inferring the positions of its nodes within this space from real data. Previous approaches typically rely on one-mode projections of bipartite networks –using, for instance, the D-Mercator tool García-Pérez et al. (2019); Jankowski et al. (2023)– to embed each type of node. However, such projections often introduce artificial correlations and inflated clustering, reducing embedding accuracy. Embeddings of one-mode projections have been used in Serrano et al. (2012) to infer the coordinates of one node type, which are then used to estimate the coordinates of the other type. Another approach Aliakbarisani et al. (2025) transfers coordinates from a unipartite layer –when such a layer exists– and assumes that the same coordinates apply in the bipartite representation. However, none of these methods provides a model-based bipartite embedding. Finally, Kitsak et al. (2017) uses the number of common neighbors as a proxy for similarity distance between nodes, but these estimates are noisy and do not support a full embedding of the network.
@@
Here, following results in Serrano et al. (2012); Kitsak et al. (2017), we introduce the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model in which nodes of both types lie in the same $D$-dimensional similarity space. Using this model, we propose B-Mercator, an algorithm designed specifically for bipartite networks that enables the creation of multidimensional hyperbolic maps of real bipartite datasets. It is important to emphasize that choosing an appropriate similarity-space dimension $D$ is crucial: while many networks are well captured with $D=1$, others require higher $D$ to prevent distinct communities from being artificially mixed Almagro et al. (2022); Jankowski et al. (2023). To illustrate its capabilities, we embedded and analyzed three datasets: Unicodelang Unicode Consortium (2024), which links countries to the languages spoken in those regions, Metabolic King et al. (2016), in which metabolites are connected through chemical reactions, and Flavor Ahn et al. (2011), which connects food ingredients to their corresponding flavor compounds. Furthermore, we show how B-Mercator can be applied to supervised machine learning tasks, including node classification and link prediction, yielding a significant performance gain with respect to state of the art methods, especially when a strong correlation exists between nodes’ labels and their feature distributions.
I Results
I.1 The bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model and B-Mercator
@@
Figure 1: Schematic representation of the bipartite-$\mathbb{S}^{D}$ model in dimension (a) $D=1$ and (b) $D=2$. Nodes A are shown as circles whereas nodes B are shown as squares whose sizes are proportional to the nodes’ expected degrees. The angular distances between nodes A and B are highlighted ($\Delta\theta_{A_{1},B_{1}}$ and $\Delta\theta_{A_{2},B_{2}}$). Light grey lines represent the edges in the bipartite network generated by Eq. (1).
@@
The bipartite-$\mathbb{S}^{D}$ model generates bipartite networks with type-A and type-B nodes by assigning them coordinates on the surface of a $D$-sphere and expected degrees. Pairs of type-A and type-B nodes are then connected with a probability that depends on their spherical distance, rescaled by the product of their expected degrees. Figure 1 shows two examples generated by the bipartite-$\mathbb{S}^{D}$ model for $D=1$ and $D=2$, where node size represents expected degree. The model has a purely geometric counterpart, the bipartite-$\mathbb{H}^{D+1}$ model, in which expected degrees are mapped to a radial coordinate so that nodes are placed in the $(D+1)$-sphere in the hyperbolic space and the connection probability becomes a function of the hyperbolic distance only. Thus, in the hyperbolic representation, angular coordinates on the $D$-sphere encode similarity, whereas the radial coordinate encodes popularity as reflected by the expected degree. In both representations, the model includes an inverse temperature $\beta_{b}$ that modulates noise and therefore controls the coupling between network topology and geometry.
Given a real bipartite network, we treat it as a likely realization of the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model. B-Mercator then finds an embedding of both node types on the surface of a $D$-sphere $\mathbb{S}^{D}$ and their expected degrees (or, equivalently, in $\mathbb{H}^{D+1}$) that maximizes the likelihood under the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model. A full description of the model and the technical details of the B-Mercator embedding algorithm are provided in Methods, Sections III.1 and III.2.
I.2 Validation
@@
To validate our method, we generated single realizations of synthetic bipartite networks from the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model with different dimensions and parameters. We then used B-Mercator for each such synthetic network to infer both the model parameters and node coordinates $(\kappa^{\text{inf}},\theta^{\text{inf}})$, and compared them with the true values $(\kappa^{\text{true}},\theta^{\text{true}})$. Figure 2 shows comparisons between the true and inferred coordinates for type-A and -B nodes. One can observe a high Spearman correlation coefficient for $D=1$ and $D=2$, corroborating the effectiveness of our embedding technique. For more examples in $D=1$ and $D=2$ as well as $D=3$, see Supplementary Figures 2-4. B-Mercator can also infer the inverse temperature $\beta_{b}$ as shown in Supplementary Figure 7.
@@
Figure 2: Validation of B-Mercator on synthetic bipartite networks. Relationship between the original and the inferred coordinates of the (a) bipartite-$\mathbb{S}^{1}$ and (b) bipartite-$\mathbb{S}^{2}$ models. In the top left corner of each figure, we report the value of the Spearman correlation coefficient between the inferred and original coordinates. Since the inferred coordinates might be rotated, we transform them to minimize the average angular distance between the original and inferred coordinates (Supplementary Section IV in Jankowski et al. (2023)). With parameters: $N_{A}=500$ (number of type-A nodes), $N_{B}=1000$ (number of type-B nodes), $\gamma_{A}=2.7$ (exponent of the powerlaw degree distribution of type-A nodes), $\gamma_{B}=2.1$ (exponent of the powerlaw degree distribution of type-B nodes), $\langle k_{A}\rangle=10$ (average degree of type-A nodes), $\beta_{b}=1.5D$ (inverse temperature), with dimension $D=1$ for (a) and $D=2$ for (b).
As a further check, we tested the reproducibility of the topological properties of the original network embedded by B-Mercator. Using the model parameters and coordinates inferred by B-Mercator, we generated an ensemble of synthetic networks using the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model (see Supplementary Figures 8–12). We then evaluated ensemble averages and standard deviations for various topological properties and compared them with those of the original network. The degree distributions and clustering spectra of type-A and type-B nodes were very well reproduced. We also observed good agreement between the empirical and theoretical connection probabilities. These findings confirm that B-Mercator accurately reconstructs the nodes’ coordinates in synthetic networks and reliably determines other model parameters, such as hidden degrees and inverse temperature, regardless of the network dimensionality.
I.3 Bipartite greedy routing
@@
In order to establish a meaningful geometric representation of a bipartite network at the global scale, we introduce bipartite greedy routing (BGR) as a practical tool to infer the network’s effective dimension Boguñá et al. (2009). The idea is to first embed the bipartite network into a latent geometric space with B-Mercator and then test how well nodes can route information by simply forwarding messages to their neighbors closest to the destination in that space. By systematically evaluating the success of these greedy routes –measured, for instance, by the probability that messages reach their targets without getting stuck– we gain insights into the dimensional structure underlying the network. In this way, the performance of greedy routing serves as an indicator of how well the bipartite graph can be embedded in a space of a given dimension, effectively allowing us to determine the dimension that best captures its structure.
@@
Figure 3: Bipartite greedy routing in synthetic networks. (a) Schematic view of the greedy routing protocol. We select a type-A node as an origin ($\mathbf{S_{A}}$) and a type-B ($\mathbf{T_{B}}$) as a destination. The red arrows show how the message is forwarded towards the destination. In the second example, we select two type-B nodes as source ($\mathbf{S_{B^{\prime}}}$) and destination ($\mathbf{T_{B^{\prime\prime}}}$) and outline the greedy path with purple color. The line width is proportional to the connection probability (Eq. 1). (b) Success rate as a function of embedded dimension for networks generated in $D=\{1,2,3,4\}$. We consider here a navigation protocol where a source is a type-A node and a destination a type-B one. Results are obtained by averaging over 10 realizations with parameters: $N_{A}=500$ (number of type-A nodes), $N_{B}=500$ (number of type-B nodes), $\gamma_{A}=2.5$ (exponent of the powerlaw degree distribution of type-A nodes), $\gamma_{B}=2.5$ (exponent of the powerlaw degree distribution of type-B nodes), $\langle k_{A}\rangle=10$ (average degree of type-A nodes), $\beta_{b}=1.5D$ (inverse temperature). The box ranges from the first quartile to the third quartile. A horizontal line goes through the box at the median. The whiskers go from each quartile to the minimum or maximum.
We implemented a BGR protocol in which both the origin and destination can be either type-A or type-B nodes thus defining four variants. In Figure 3a, we depict a schematic picture of the BGR for the A–B variant, i.e., when the source node is a type-A node ($S_{A}$), and the target node is a type-B node ($T_{B}$). The message is forwarded from $S_{A}$ to the type-B node that is hyperbolically closest to the target. Since the node $B_{1}$ is not the destination, the message is forwarded again to a type-A node. The process is repeated until the destination is reached or the message becomes stuck. Then, the BGR protocol is executed for a large number of randomly chosen node pairs to assess the global network’s geometric properties.
We tested the BGR protocol in synthetic networks generated from the bipartite-$\mathbb{S}^{D}$ model. First, we generated networks with specific dimensionality and topological properties, and we obtained their hyperbolic maps by embedding them using B-Mercator with different embedding dimensions. The performance of BGR was assessed based on two key measures: the proportion of messages that successfully reach their destination $p_{s}$, i.e., the success rate, and the mean stretch, where stretch is defined for each path connecting a source and target node as the ratio of the hop count of a successful greedy path to the shortest path. In Figure 3b, we show the success rate as a function of the embedding dimension for networks generated using the bipartite-$\mathbb{S}^{D}$ model with dimensions ranging from $D=1$ to $D=4$. The performance of BGR is optimal (in terms of $p_{s}$ and mean stretch, see Supplementary Figure 14) when B-Mercator is used with the same dimension that was used to generate the network, thus justifying BGR as an alternative method to infer the effective dimension of real networks, different from the topological-based method introduced in Almagro et al. (2022). These results are consistent across all variants of BGR (see Supplementary Figure 13) and corroborate the results obtained for unipartite networks Jankowski et al. (2023).
I.4 Embedding of real bipartite networks